Question 10 (4 points)

 

On the + y axis a laser is located at y = + 3.0 cm. The coordinates of a small target are x = + 12.0 cm and y = + 6.0 cm. The + x axis represents the edge-on view of a plane mirror. At what point on the + x axis should the laser be aimed in order for the laser light to hit the target after reflection?

Question 10 options:


A) 

+4 cm


B) 

+3 cm


C) 

-3 cm


D) 

- 4 cm



1
Expert's answer
2021-11-14T17:13:01-0500

Angle of reflection is the angle of incidence. Thus, angles between the rays and the x-axis must be equal:


"\frac{y_1}{x_1}=\frac{y_2}{x_2-x_1},\\\space\\
\frac3{x_1}=\frac6{12-x_1},\\
x_1=4\text{ cm}."